OpenStudy (anonymous):

"a metal stays solid at fahrenheit temperatures below 1553.1. Dtermine in terms of an inequality those celsius temps for which the metal stays solid." use the formula f=9/5c+32

OpenStudy (anonymous):

Solve the formula for c, and since you want the temp to be belwo a certain value replace the + with a < symbol\[c<\frac{5}{9}\left( f-32 \right)\]and this is your answer. To find the actual temp in celcsius just sub 1553.1 for f in the above equation\[c<\frac{5}{9}(1553.1-32)=845.1\]